OpenStudy (anonymous):

An ellipse has foci (0, ±1) and vertices (0, ±2). What is the eccentricity of the ellipse? 0.5 right?

jimthompson5910 (jim_thompson5910):

I'm getting the same thing based on this http://www.mathopenref.com/ellipseeccentricity.html so e = c/a = 1/2 = 0.5